\begin{thebibliography}{999}


\bibitem{ak}
Albiac, Fernando; Kalton, Nigel J.
Topics in Banach space theory.
Graduate Texts in Mathematics, 233. Springer, New York, 2006.

\bibitem{aat}
 Argyros, Spiros A.; Arvanitakis, Alexander D.; Tolias, Andreas G.
 Saturated extensions, the attractors method and hereditarily James tree spaces.
 Methods in Banach space theory, 1Ð90, London Math. Soc. Lecture Note Ser., 337, Cambridge Univ. Press, Cambridge, 2006.


 \bibitem{bgiks} Berinde, R.; Gilbert, A.~C.; Indyk, P; Karloff, H.; Strauss, M.~J.
Combining geometry and combinatorics: a unified approach to sparse signal recovery.
2008 46th Annual Allerton Conference on Communication, Control, and Computing (2008), 798--805.


  \bibitem{djt}
  Diestel, Joe; Jarchow, Hans; Tonge, Andrew. Absolutely summing operators. Cambridge Studies in Advanced Mathematics, 43. Cambridge University Press, Cambridge, 1995.

  \bibitem{gon}
Gonz\'alez, Manuel; Mart\'inez-Abej\'on, Antonio. Tauberian operators. Operator Theory: Advances and Applications, 194. BirkhŠuser Verlag, Basel, 2010.

  \bibitem{go}
Gonz\'alez, Manuel; Onieva, Victor M. On the instability of non-semi-Fredholm operators under compact perturbations. J. Math. Anal. Appl. 114 (1986), no. 2, 450--457.


 \bibitem{lt2}
Lindenstrauss, Joram; Tzafriri, Lior
Classical Banach spaces. II. Function spaces.
Ergebnisse der Mathematik und ihrer Grenzgebiete, 97. Springer-Verlag, Berlin-New York, 1979.

 \bibitem{mac}
Mackey, George W.
Note on a theorem of Murray.
Bull. Amer. Math. Soc. 52, (1946),  322--325.




\bibitem{A.N.} A. B. Nasseri, \textit{The spectrum of operators on $C(K)$ with the Grothendieck property and characterization of $J$-Class Operators which are adjoints.}



  \bibitem{bah} Nasseri, Amir Bahman.
http://mathoverflow.net/questions/101253

  \bibitem{roy}
Royden, H. L.
Real analysis.
Third edition.
Macmillan Publishing Company, New York, 1988.




 \end{thebibliography}

\documentclass[12pt]{article}
\usepackage{amsfonts,amssymb,epsfig}
\usepackage{amsmath}

\date{}
\begin{document}
\newtheorem{df}{Definition}
\newtheorem{thm}{Theorem}

\newtheorem{lm}{Lemma}
\newtheorem{pr}{Proposition}
\newtheorem{co}{Corollary}
\newtheorem{re}{Remark}
\newtheorem{note}{Note}
\newtheorem{claim}{Claim}
\newtheorem{problem}{Problem}

\def\R{{\mathbb R}}

\def\E{\mathbb{E}}
\def\calF{{\cal F}}
\def\N{\mathbb{N}}
\def\calN{{\cal N}}
\def\calH{{\cal H}}
\def\n{\nu}
\def\a{\alpha}
\def\d{\delta}
\def\t{\theta}
\def\e{\varepsilon}
\def\t{\theta}
\def\g{\gamma}
\def\G{\Gamma}
\def\b{\beta}
\def\pf{ \noindent {\bf Proof: \  }}

\newcommand{\qed}{\hfill\vrule height6pt
width6pt depth0pt}
\def\endpf{\qed \medskip} \def\colon{{:}\;}
\setcounter{footnote}{0}

\renewcommand{\qed}{\hfill\vrule height6pt  width6pt depth0pt}

\title{Injective Tauberian operators on $L_1$ and operators with dense range on $\ell_\infty$
\thanks {AMS subject classification: 46E30, 46B08, 47A53 \ \
Key words: $L_1$, Tauberian operator, $\ell_\infty$}}

\author{William B. Johnson\thanks{Supported in part by NSF DMS-1301604  and U.S.-Israel Binational Science
Foundation
 }, Amir Bahman Nasseri,\\   Gideon Schechtman\thanks{Supported in part by U.S.-Israel Binational Science Foundation. Participant
 NSF Workshop in Analysis and Probability, Texas A\&M University
 }, \ and Tomasz Tkocz\thanks{T. Tkocz thanks his PhD supervisor, Keith Ball, for his invaluable constant
advice and encouragement}} \maketitle

\begin{abstract}
There exist injective Tauberian operators on $L_1(0,1)$ that have dense, non closed range. This gives injective, non surjective operators on $\ell_\infty$ that have dense range. Consequently, there are two quasi-complementary, non complementary  subspaces of $\ell_\infty$ that are isometric to $\ell_\infty$.

\end{abstract}

\section{Introduction}

A (bounded, linear) operator $T$ from a Banach space $X$ into a Banach space $Y$ is called Tauberian provided $T^{**-1}Y=X$.  The structure of Tauberian operators when the domain is an $L_1$ space is well understood and exposed in Gonz\'ales and Mart\'inez-Abej\'on's book \cite[Chapter 4]{gon}. (For convenience they only consider $L_1(\mu)$ when $\mu$ is finite and purely nonatomic, but their proofs for the results we mention work for general $L_1$ spaces.) In particular,
\cite[Theorem 4.1.3]{gon}    implies that when $X$ is an $L_1$ space,   an  operator $T:X\to Y$ is Tauberian iff whenever $(x_n)$ is a sequence of disjoint unit vectors, there is an $N$ so that the restriction of $T$ to $[x_N]_{n=N}^\infty$ is an isomorphism (and, moreover, the norm of the inverse of the restricted operator is bounded independently of the disjoint sequence). From this it follows that an injective operator $T:X\to Y$ is Tauberian iff it isomorphically preserves isometric copies of $\ell_1$ in the sense that  the restriction of $T$  to any subspace of $X$ that is isometrically isomorphic to $\ell_1$ is an isomorphism. (Recall that a subspace of an $L_1$ space is isometrically isomorphic to $\ell_1$ iff it is the closed linear span of a sequence of non zero disjoint vectors  \cite[Chapter 14.5]{roy}.) Since $Tu$ is Tauberian if $T$ is Tauberian and $u$ is an isomorphism, one deduces that an injective Tauberian operator from an $L_1$ space isomorphically preserves isomorphic copies of $\ell_1$ in the sense that  the restriction of $T$ to any subspace of $X$ that is isomorphic to $\ell_1$ is an isomorphism. Thus injective Tauberian operators from an $L_1$ space are opposite to $\ell_1$-singular operators; i.e., operators whose restriction to every  subspace isomorphic to $\ell_1$ is {\em not} an isomorphism.

The main result in this paper is a negative solution to \cite[Problem 1]{gon}: Suppose $T$ is a Tauberian operator on an $L_1$ space.  Must $T$ be upper semi-Fredholm; i.e., must the range $\mathcal{R}(T)$  of $T$ be closed and the null space $\mathcal{N}(T)$ of $T$ be finite dimensional?  The basic example is a Tauberian operator on $L_1(0,1)$ that has infinite dimensional null space.  This is rather striking because the Tauberian condition is equivalent to the statement that there is $c>0$ so that the restriction of the operator to $L_1(A)$ is an isomorphism whenever the subset $A$ of $[0,1]$ has Lebesgue measure at most $c$.

In fact, we show that there is an injective, dense range, non surjective Tauberian operator on $L_1(0,1)$.  Since $T$ is Tauberian, $T^{**}$ is also injective, so  $\mathcal{R}(T^*)$ is dense and proper, and  $T^*$ is injective because  $\mathcal{R}(T)$ is dense. This solves a problem \cite{bah} the second author raised on MathOverFlow.net that led to the collaboration of the authors.

\section{The examples}

We begin with a lemma that is an easy consequence of characterizations of Tauberian operators on $L_1$ spaces.

\begin{lm}\label{finite} Let $X$ be an $L_1$ space and  $T$ an operator from $X$ to a Banach space $Y$.  The operator $T$ is Tauberian if and only if there is $r>0$ and a natural number $N$ so that if $(x_n)_{n=1}^N$ are disjoint unit vectors in $X$, then $\max_{1\le n \le N} \|Tx_n\| \ge r$.
\end{lm}
\pf The condition in the lemma clearly implies that if $(x_n)$ is a disjoint sequence of unit vectors in $X$, then $\liminf_n \|Tx_n\| >0$, which is one of the equivalent conditions for $T$ to be Tauberian \cite[Theorem 4.1.3]{gon}.  On the other hand, suppose that there are disjoint collections $(x^n_k)_{k=1}^n$, $n=1,2,\dots$ with $ \max_{1\le k \le n } \|T{x^n_k} \| \to 0$ as $n \to \infty$.  Then the closed sublattice generated by $\cup_{n=1}^\infty (x^n_k)_{k=1}^n$ is a separable abstract $L_1$ space (meaning that it is a Banach lattice such that $\|x+y\|= \|x\|+\|y\|$ whenever $|x| \vee  |y| =0$)  and hence is  order isometric to $L_1(\mu)$ for some probability $\mu$ by Kakutani's theorem (see e.g. \cite[Theorem 1.b.2]{lt2}). Choose $1\le k(n) \le n$ so that the support of $x_{k(n)}^n$ in $L_1(\mu)$ has measure at most $1/n$.  Since $T$ is Tauberian, by \cite[Proposition 4.1.8]{gon} necessarily $\liminf_n \|Tx_{k(n)}^n \| >0$, which is a contradiction. \qed

\medskip
The reason that Lemma \ref{finite} is useful for us is that the condition in the Lemma is stable under ultraproducts.  Call an operator that satisfies the condition in Lemma \ref{finite} $(r,N)$-Tauberian.  For background on ultraproducts   of Banach spaces and of operators, see \cite[Chapter 8]{djt}.  We use the fact that the ultraproduct of $L_1$ spaces is an abstract $L_1$ space and hence   is order isometric to $L_1(\mu)$ for some measure $\mu$.

\begin{lm}\label{ultra} Let $(X_k)$ be a sequence of $L_1$ spaces, and for each $k$  let $T_k$  be a norm one linear operator from $X_k$ into a Banach space $Y_k$.  Assume that   there is $r>0$ and a natural number $N$ so that each operator $T_k$ is  $(r,N)$-Tauberian. Let $\mathcal{U}$ be a free ultrafilter on the natural numbers.  Then $(T_k)_\mathcal{U}: (X_k)_\mathcal{U} \to (Y_k)_\mathcal{U}$
is $(r,N)$-Tauberian.
\end{lm}
Here $(T_k)_\mathcal{U}$ is the usual ultraproduct of the sequence $(T_k)$, defined by  \hfill\break

$(T_k)_\mathcal{U}(x_k) = (T_kx_k)$.

\vskip9pt
\pf  The vectors $(x_k)$ and $(y_k)$ are disjoint in the abstract $L_1$ space $(X_k)_\mathcal{U}$ iff $\lim_\mathcal{U} \||x_k|\wedge |y_k|\| =0$, so it is only a matter of proving that if $T$ is $(r,N)$-Tauberian from some $L_1$ space $X$, then for each $\e>0$ there is $\d>0$ so that if $x_1,\dots, x_N$ are unit vectors in $X$ and $\| |x_n|\wedge |x_m| \| < \d $ for $1\le n<m \le N$, then $\max_{1\le n \le N }\|Tx_n\| > r-\e$. But if $x_1,\dots, x_N$ are unit vectors that are $\e$-disjoint as above, and $y_1,\dots,y_n$ are defined by \hfill\break

$y_n := [|x_n|-(|x_n|\wedge (\vee \{|x_m| : m\not= n\})]\text{sign}(x_n)$, \hfill\break

\noindent then the $y_n$ are disjoint and all have norm at least $1-N\d$.  Normalize the $y_n$ and apply the $(r,N)$-Tauberian condition to this normalized disjoint sequence to see that $\max_{1\le n \le N }\|Tx_n\| > r-\e$ if $\d = \d(\e, N)$ is sufficiently small. \qed

\medskip
An example that answers   \cite[Problem 1]{gon} is the restriction of an ultraproduct of operators on finite dimensional $L_1$ spaces constructed in \cite{bgiks}.

\begin{thm}\label{theorem1}
There is a Tauberian operator $T$ on $L_1(0,1)$ that has an infinite dimensional null space.  Consequently, $T$ is not upper semi-Fredholm.
\end{thm}
\pf An immediate consequence of \cite[Proposition 6 \& Theorem 1]{bgiks} is that there is $r>0$ and a natural number $N$ so that for all sufficiently large $n$ there is a norm one   $(r,N)$-Tauberian operator $T_n$ from $\ell_1^n$ into itself with $\dim \mathcal{N}(T_n) > rn$. The ultraproduct $\tilde{T}:=(T_n)_\mathcal{U}$ is then a norm one
$(r,N)$-Tauberian operator on the gigantic $L_1$ space $X_1:=(\ell_1^n)_\mathcal{U}$, and the null space of $\tilde{T}$ is infinite dimensional.  Take any separable infinite dimensional subspace $X_0$ of $\mathcal{N}(\tilde{T})$ and let $X$ be the closed sublattice of $X_1$   generated by $X_0$.  Let $Y$ be the sublattice of $X_1$ generated by $\tilde{T}X$ and let $T$ be the restriction of $\tilde{T}$ to $X$, considered as an operator into $Y$.  So $X$ and $Y$ are separable $L_1$ spaces and by Lemmas \ref{finite} and \ref{ultra} the operator $T$ is Tauberian.  Of course, by construction $\mathcal{N}(T)$ is infinite dimensional and reflexive (because $T$ is Tauberian).  Thus $X$ is not isomorphic to $\ell_1$ and hence is isomorphic to $L_1(0,1)$.  So is $Y$, but that does not matter: $Y$, being a separable $L_1$ space, embeds isometrically into $L_1(0,1)$. \qed

We want to ``soup up" the operator $T$ in Theorem \ref{theorem1} to get an injective, non surjective, dense range Tauberian operator on $L_1(0,1)$.  We could quote a general result \cite[Theorem 3.4]{go} of Gonz\'alez and Onieva to shorten the presentation, but we prefer to give a short direct proof.

We recall a simple known lemma:

\begin{lm}\label{nuclear}
Let $X$ and $Y$ be separable infinite dimensional Banach spaces and $\e>0$. Let $Y_0$ be a countable dimensional dense subspace of $Y$.Then there is a nuclear operator $u:X\to Y$ so that $u$ is injective and $\|u\|_\wedge < \e$ and $uX \supset Y_0$.
\end{lm}
\pf Recall that an $M$-basis for a Banach space $X$  is a biorthogonal system $(x_\a,x^*_\a)\subset X\times X^*$ such that the linear span of $(x_\a)$ is dense in $X$ and $\cap_\a \mathcal{N}(x^*_\a) =\{0\}$.  Every separable Banach space $X$ has an $M$-basis \cite{mac}; moreover, the vectors $(x_\a)$  in the $M$-basis can span any given countable dimensional dense subspace of $X$.

Take $M$-bases $(x_n,x^*_n)$ and $(y_n,y^*_n)$ for $X$ and $Y$, respectively, normalized so that $\|x_n^*\|=1=\|y_n\|$ and such that the linear span of $(y_n)$ is $Y_0$. Choose $\lambda_n>0$ so that $\sum_n \lambda_n < \e$ and set $u(x)= \sum_n \lambda_n \langle x^*_n, x \rangle y_n$. \qed

\begin{thm}\label{theorem2}
There is an injective, non surjective, dense range Tauberian operator on $L_1(0,1)$.
\end{thm}
\pf By Theorem \ref{theorem1} there is a  Tauberian operator $T$ on $L_1(0,1)$ that has an infinite dimensional null space.  By Lemma \ref{nuclear} there is a nuclear operator $\tilde{v} : \mathcal{N}(T) \to L_1(0,1)$ that is injective and has dense range, and we can extend $\tilde{v} $  to a nuclear operator $v$  on $L_1(0,1)$. We can choose $\tilde{v}$ so that $\tilde{v}(\mathcal{N}(T) ) \cap TL_1(0,1)$ is infinite dimensional by the last  statement in Lemma \ref{nuclear}.  This guarantees that the Tauberian operator $T_1:= T+v$ has an infinite dimensional null space (this allows us to avoid breaking the following argument into cases).

Now $\mathcal{N}(T_1) \cap \mathcal{N}(T) = \{0\}$, so again by Lemma \ref{nuclear} and the extension property of nuclear operators there is a nuclear operator $u : L_1(0,1)/{\mathcal{N}(T)} \to \ell_1$ so that the restriction of $u$ to $ Q_{\mathcal{N}(T)} \mathcal{N}(T_1)$ is injective and has dense range (here for a subspace $E$ of $X$, the operator $Q_E$ is the quotient mapping from $X$ onto $X/E$). Finally, define $T_2 : L_1(0,1) \to L_1(0,1) \oplus_1 \ell_1$ by $T_2 x := T_1x \oplus u Q_{\mathcal{N}(T)} x$.  Then $T_2$ is an injective Tauberian operator with dense range.  $T_2$ is not surjective because $P_{\ell_1} T_2$ is nuclear by construction, where $P_{\ell_1}$ is the   projection of $L_1(0,1) \oplus_1 \ell_1$ onto $\{0\} \oplus_1 \ell_1$.  Since $L_1(0,1) \oplus_1 \ell_1$ is isomorphic to $L_1(0,1)$, this completes the proof. \qed

\begin{co} \label{corollary} There is an injective, dense range, non surjective operator on $\ell_\infty$.  Consequently, there is a quasi-complementary, non complementary decomposition of $\ell_\infty$ into two subspaces each of which is  isometrically isomorphic to $\ell_\infty$.
\end{co}
\pf Let $T$ be an injective, dense range, non surjective Tauberian operator on $L_1(0,1)$ (Theorem \ref{theorem2}).  Since $T$ is Tauberian, $T^{**}$ is also injective, so $T^*$ has dense  range but $T^*$ is not surjective because its range is not closed, and  $T^*$ is injective because $T$ has dense range.  The operator $T^*$ translates to an operator on $\ell_\infty$ that has the same properties because $L_\infty$ is isomorphic to $\ell_\infty$ by an old result due to Pe\l czy\'nski (see, e.g., \cite[Theorem 4.3.10]{ak}) (notice however that, unlike $T^*$,  the operator on $\ell_\infty$ cannot be weak$^*$ continuous).

For the ``consequently" statement, let $S$ be any norm one  injective, dense range, non surjective operator on $\ell_\infty$.  In the space $\ell_\infty \oplus_\infty \ell_\infty$, which is isometric to $\ell_\infty$, define
$X:= \ell_\infty \oplus \{0\}$ and $Y:= \{(x,Sx) : x\in \ell_\infty\}$. Obviously $X$ and $Y$ are isometric to $\ell_\infty$ and $X+Y= \ell_\infty \oplus S \ell_\infty$, which is a dense proper subspace of  $\ell_\infty \oplus_\infty \ell_\infty$.  Finally, $X\cap Y =\{0\}$ since $S$ is injective, so $X$ and $Y$ are quasi-complementary, non complementary subspaces of $\ell_\infty \oplus_\infty \ell_\infty$. \qed

Theorem \ref{theorem2} and the MathOverFlow question \cite{bah} suggest the following problem: Suppose $X$ is a separable Banach space (so that $X^*$ is isometric to a weak$^*$ closed subspace of $\ell_\infty$) and $X^*$ is non separable.  Is there a dense range operator on $X^*$ that is not surjective?  The answer is ``no":  Argyros, Arvanitakis, and Tolias \cite{aat} constructed a separable space $X$ so that $X^*$ is non separable, hereditarily indecomposable (HI), and every strictly singular operator on $X^*$ is weakly compact. Since $X^*$ is HI, every operator on $X^*$ is of the form $\lambda I +S$ with $S$ strictly singular.  If $\lambda \not= 0$, then  $\lambda I +S$  is Fredholm of index zero by Kato's classical perturbation theory. On the other hand, since every weakly compact  subset of the dual to a separable space is norm separable, every  strictly singular operator on $X^*$ has separable range.  (Thanks to Spiros Argyros for bringing this example to our attention.)

Any operator $T$ on $l^\infty$ that has dense range but is not surjective has the property that $0$ is an interior point of $\sigma(T)$. This follows from Thm 2.6 in \cite{A.N.}, where it is shown that $\partial\sigma(T)\subset\sigma_{p}(T^{*})$ for any operator $T$ acting on a $C(K)$ space which has the Grothendieck property.


 \begin{tabular}{l}
W.~B.~Johnson\\
Department of Mathematics\\
Texas A\&M University\\
College Station, TX  77843 U.S.A.\\
{\tt johnson@math.tamu.edu}
\\
\end{tabular}

\bigskip

  \begin{tabular}{l}
A.~B.~Nasseri\\
Fakult\"at f\"ur Mathematik\\
Technische Universit\"at Dortmund\\
 D-44221 Dortmund, Germany\\
{\tt amirbahman@hotmail.de}
\\
\end{tabular}

\bigskip

\begin{tabular}{l}
G.~Schechtman\\
Department of Mathematics\\
Weizmann Institute of Science\\
Rehovot, Israel\\
{\tt gideon@weizmann.ac.il}\\
\end{tabular}

\bigskip

\begin{tabular}{l}
T.~Tkocz\\
Mathematics Institute\\
University of Warwick\\
Coventry CV4 7AL, UK\\
{\tt ttkocz@gmail.com}
\end{tabular}

\end{document}